# Advances in Protein Representation Learning: Methods, Applications, and Future Directions

### v1

@@ Abstract
Table 2: Summary of widely used databases in Protein Representation Learning (PRL), categorized by sequence, structure, and function.
Name Data Description
UniProtKB 102, 103 Sequence, Function UniProtKB comprises two main components: Swiss-Prot, which contains manually curated protein sequences, and TrEMBL, which provides automatically annotated sequences from high-throughput studies.
UniRef 104 Sequence UniRef is a clustering system based on the UniProt protein database, reducing sequence redundancy and improving computational efficiency by grouping sequences at different identity thresholds.
Pfam 105 Sequence Pfam is a database of protein families and domains, where each family is defined by multiple sequence alignments and profile Hidden Markov Models (HMMs) to facilitate functional annotation.
@@
MGnify 106 Sequence MGnify is a large-scale metagenomic database containing over 2.4 billion non-redundant protein sequences predicted from environmental microbiomes.
@@
PDB 107 Structure The Protein Data Bank (PDB) is a repository of experimentally determined biomolecular structures, obtained through techniques such as X-ray crystallography, NMR spectroscopy, and cryo-electron microscopy.
@@
AlphaFold DB 108 Structure AlphaFold DB stores protein structures predicted using the AlphaFold series, offering high-confidence computational models that expand structural coverage beyond experimentally resolved proteins.
ESMAtlas 108 Structure ESMAtlas is a large-scale protein structure prediction database generated using ESMFold, containing over 617 million predicted structures, including millions of novel proteins not yet experimentally characterized.
@@
Gene Ontology 109 Function The Gene Ontology (GO) database provides structured annotations of protein functions, categorizing them into biological processes, molecular functions, and cellular components to support functional characterization.
8 8. Applications
@@
Protein Representation Learning (PRL) has revolutionized biological research and biomedical innovation, providing data-driven solutions for understanding and engineering proteins. By learning meaningful representations of protein sequences, structures, and interactions, PRL enables a wide range of applications, from predicting fundamental protein properties to advancing drug discovery. As illustrated in Fig. 4, the applications of PRL can be broadly categorized into four main domains: protein property prediction, protein structure prediction, protein design and optimization, and drug discovery.
@@
Figure 4: Overview of key applications of Protein Representation Learning (PRL). PRL enables advancements in multiple domains, including protein property prediction, structure prediction, protein design and optimization, and drug discovery.
8.1 8.1. Protein Property Prediction
@@
Benchmarks for evaluating protein property prediction models are categorized based on the type of property being assessed. For individual protein properties, Tasks Assessing Protein Embeddings (TAPE) 110 provides a standardized evaluation framework, including fluorescence prediction and protease stability assessment. Another widely used benchmark, UniProt 102, serves as a comprehensive database of protein sequences, functional annotations, and metadata from diverse organisms, offering a valuable resource for assessing sequence-based property predictions across various protein families and functional contexts. In contrast, interaction-based properties—such as protein-ligand and protein-protein interactions—require different benchmarks that incorporate structural and binding information. Prominent datasets in this category include PDBBind 111, Leak Proof PDBBind 112, Davis 113, KIBA 114, PLINDER 115 and PINDER 116, which evaluate models based on binding affinity predictions. These datasets serve as widely recognized standards, enabling consistent model comparison and performance assessment across different protein property prediction tasks.
@@
With the remarkable advancements in Protein Language Models (PLMs), most current research leverages these models to extract meaningful representations of protein sequences, applying them to a wide range of property prediction tasks. These models have been successfully employed for both individual protein properties 7, 117, 118 as well as interaction-based properties 119, 93, 120. Beyond sequence-based representations, many efforts have explored integrating multiple modalities, such as structural, physicochemical, and functional data, to further enhance protein representations beyond what PLMs alone can achieve 58, 42, 72, 121. By incorporating additional biological information, these approaches aim to provide a more comprehensive view of protein function and interactions, particularly in tasks where spatial and contextual features play a significant role, such as protein-ligand binding and protein-protein interactions.
8.2 8.2. Protein Structure Prediction
@@
Protein structure prediction, also known as protein folding, aims to determine a protein’s three-dimensional conformation from its amino acid sequence. Since protein structure dictates function, accurately predicting spatial arrangements is essential for understanding molecular interactions, drug binding, and enzymatic activity. Traditionally, protein structures have been determined using experimental techniques such as X-ray crystallography 122 and cryo-electron microscopy (Cryo-EM) 123, which provide high-resolution structural insights. However, these methods are often time-consuming, costly, and limited by challenges such as crystallization feasibility and sample preparation 124. Advances in computational structure prediction have transformed the field by providing faster, scalable alternatives that complement experimental approaches, broadening the scope of protein engineering and biomedical research.
@@
To evaluate the accuracy of structure prediction models, several benchmarking platforms have been established. The Critical Assessment of Techniques for Protein Structure Prediction (CASP) 125, a biennial blind assessment, is widely regarded as the gold standard for evaluating protein folding models. In addition, the Continuous Automated Model EvaluatiOn (CAMEO) 126 provides real-time assessments by continuously comparing predictions against newly released experimental structures. These benchmarks are instrumental in validating and refining computational models, ensuring their reliability and applicability in structural biology.
@@
Protein structure prediction models can be broadly categorized into MSA-based and MSA-free approaches. MSA-based methods leverage multiple sequence alignments to capture evolutionary relationships, often resulting in higher accuracy for proteins with rich evolutionary data. Notable models in this category include AlphaFold 8, RosettaFold 127, OpenFold 128, and MSAGPT 11. In contrast, MSA-free methods predict structures directly from single sequences without relying on evolutionary context, making them more efficient and applicable to proteins with limited sequence homology. Examples of MSA-free models include ESMFold 13, OmegaFold 129, and HelixFold-Single 130. While these models have greatly improved protein structure prediction, they are primarily limited to static conformations, as they are trained on crystallographic and experimentally resolved structures. However, proteins are inherently dynamic, undergoing conformational changes essential to their function. To address this limitation, recent advancements have introduced generative models capable of predicting conformational ensembles, such as BioEmu 131.
8.3 8.3. Protein Design and Optimization
@@
Protein design and optimization involve strategies for engineering proteins with specific structural and functional properties. Broadly, these strategies fall into two main categories: sequence-based design, which optimizes amino acid sequences to achieve target properties, and structure-based design, which leverages protein structural features—either by generating backbones and applying inverse folding to identify compatible sequences or by co-designing sequence and structure to enhance functional and structural outcomes. These approaches have been applied to various protein types, including individual proteins with optimized functionality, ligand-binding proteins for molecular interactions, antibodies for therapeutic applications, and enzymes for improved catalytic efficiency in industrial and biomedical settings.
8.3.1 8.3.1. Individual Protein Property Optimization
@@
Individual protein property optimization focuses on enhancing a protein’s intrinsic characteristics, such as stability, activity, specificity, and catalytic efficiency, without explicitly considering interactions with other molecules. Traditional approaches, such as directed evolution with random mutagenesis 132, 133, 134, iteratively introduce mutations and select for improved variants. While these methods have been widely used and have led to many successful optimizations, they often require extensive experimental screening. Computational strategies have been developed to facilitate a more systematic exploration of sequence space, offering alternatives that reduce experimental reliance while aiming to improve efficiency in protein engineering.
@@
To evaluate optimization strategies, several benchmark datasets have been established for sequence-based optimization, each representing distinct challenges across protein targets, originating organisms, sequence lengths, dataset sizes, and fitness objectives. Notable benchmarks include Green Fluorescent Protein (avGFP) 135, Adeno-Associated Viruses (AAV) 136, TEM-1 $\beta$-Lactamase (TEM) 137, Ubiquitination Factor Ube4b (E4B) 138, Aliphatic Amide Hydrolase (AMIE) 139, Levoglucosan Kinase (LGK) 140, Poly(A)-binding Protein (Pab1) 141, and SUMO E2 Conjugase (UBE2I) 142. These datasets enable standardized assessment of different protein optimization methods.
@@
Earlier approaches focused on optimizing protein in discrete sequence space. These methods explored mutation-based search strategies, reinforcement learning, and PLMs to propose beneficial sequence modifications 143, 144. Building upon this foundation, recent advancements in PRL have introduced computational frameworks capable of mapping protein sequences into fitness landscapes. These models aim to capture sequence-function relationships and facilitate direct sequence optimization based on learned representations. Unlike experimental screening, PRL-based approaches rely on large-scale protein datasets to infer meaningful sequence features and predict fitness variations. A common strategy in fitness landscape modeling integrates sequence and fitness information within a machine-learning framework. One such approach involves jointly training an autoencoder with a prediction network, where an encoder extracts sequence features, a decoder reconstructs sequences, and a separate model predicts fitness values from the latent space 145, 146. Protein Language Models (PLMs) are often used as encoders to capture contextual sequence representations, and the prediction network guides latent space organization based on fitness variations.
@@
Other studies have explored methods to improve fitness landscape modeling through smoothing techniques, aiming to enhance search efficiency in sequence space. Graph-based smoothing techniques have been introduced to incorporate protein similarity graphs, promoting structural continuity in the fitness landscape 147, 148. These methods aim to reduce abrupt fitness changes caused by minor sequence variations, improving the robustness of optimization algorithms. Overall, while experimental and computational strategies each offer advantages and limitations, ongoing research continues to explore ways to enhance protein optimization efficiency while ensuring the reliability and interpretability of computational predictions.
8.3.2 8.3.2. Ligand-Binding Protein Design
@@
Ligand-binding protein design involves engineering proteins to interact with small molecules such as drugs, metabolites, or signaling compounds. These interactions are central to biological processes, making ligand-binding protein design essential for applications in biosensors, precision medicine, and enzyme-mediated synthesis. Achieving high specificity and affinity for target ligands requires a detailed understanding of binding site geometry, molecular interactions, and binding thermodynamics, which computational methods have been increasingly employed to model and optimize.
@@
Several benchmark datasets support the evaluation of ligand-binding protein design strategies by providing experimentally validated protein-ligand complexes. PDBbind 111 offers a curated collection of protein-ligand structures with binding affinity measurements, serving as a reference for affinity prediction and binding site modeling. CSAR (Community Structure Activity Resource) 149 provides high-resolution crystallographic data with a diverse range of ligand complexes, offering a structured framework for evaluating binding specificity across various protein targets. These datasets are widely used to assess structure-based and sequence-based ligand-binding design approaches.
@@
Computational strategies for ligand-binding protein design typically leverage structural insights from native protein-ligand complexes to refine side-chain interactions and backbone conformations, with the aim of enhancing binding affinity 150, 151, 152, 153. Many approaches follow a structure-first paradigm, in which protein backbones are generated first, followed by inverse folding techniques 154, 155 to identify sequences that can adopt predefined structural conformations. This prioritization of structure prediction before sequence determination allows for the design of proteins that conform to specific ligand-binding geometries and can also aid in addressing cases where binding sites are not well defined.
@@
Despite progress, challenges remain, particularly in the structural validation of designed binding modes. While computational docking has been applied to generate novel binders by modifying scaffolds and loop geometries, the accuracy and stability of predicted interactions often require high-resolution experimental confirmation. Additionally, many design strategies emphasize a limited set of hotspot residues for scaffold placement, which may restrict the exploration of diverse interaction modes, particularly for targets lacking well-defined pockets or binding clefts 156, 157. To address these limitations, alternative approaches have investigated ligand-binding protein design based on sequence information alone, bypassing the reliance on structural data 87. These models leverage Protein Language Models (PLMs) to extract functional sequence motifs associated with ligand binding, allowing for the design of proteins with desired binding properties even when structural information is unavailable or difficult to obtain.
8.3.3 8.3.3. Antibody Design
@@
Antibodies are specialized immune proteins that recognize and bind to specific antigens, such as pathogens, toxins, or diseased cells, marking them for neutralization or destruction. As central components of adaptive immunity, antibodies exhibit high specificity and strong binding affinity, enabling precise immune responses. Their function is primarily dictated by variable regions, particularly the complementarity-determining regions (CDRs), which mediate antigen recognition. Antibody design focuses on engineering antibodies with improved affinity, specificity, and stability, with broad therapeutic applications in cancer, autoimmune disorders, and infectious diseases.
@@
Benchmark datasets play a critical role in evaluating sequence-based and structure-based antibody design. OAS (Observed Antibody Space) 158 provides over 2 billion immune repertoire sequences from various immune states, species, and individuals, supporting sequence-based antibody modeling. SAbDab (Structural Antibody Database) 159 compiles annotated antibody structures from the PDB, including affinity data and sequence annotations, serving as a key resource for structure-based antibody modeling. AbBind 160 includes 1,101 antibody-antigen mutants across 32 complexes with experimentally measured binding free energy changes, allowing for the study of affinity variations upon mutation. These datasets support the evaluation and development of antibody design approaches, ensuring standardized model assessment across different strategies.
@@
Computational strategies for structure-based antibody design primarily focus on CDRH3 loops, which are highly variable and critical for antigen binding. Two main methodologies have emerged: (i) 3D backbone generation, which designs structurally realistic CDRH3 loops for stable antigen recognition 161, and (ii) binding affinity prediction, which models how sequence modifications impact structural stability and antigen binding energy 162. While structure-based approaches have been widely used, a major challenge is the limited availability of high-resolution antibody structures, which constrains direct structural modeling. To address this, sequence-based models leverage Protein Language Models (PLMs) trained on large-scale immune repertoires to learn patterns associated with high-affinity antibodies. These methods enable de novo antibody generation and affinity maturation without requiring structural information 163. Additionally, sequence-based antibody design can be framed as an optimization problem, where binding affinity serves as the fitness metric, guiding the search for improved variants—similar to optimization strategies used for individual protein properties 164, 165. A promising direction in antibody design is the integration of sequence and structural information, leading to sequence-structure co-design approaches 166, 167, 168, 169, 170. By combining both modalities, these models aim to capture relationships between sequence variability and structural adaptation, facilitating the precise engineering of antibody specificity, stability, and affinity. This hybrid framework expands the design space beyond naturally occurring immune repertoires, potentially accelerating antibody discovery and improving antigen recognition strategies.
8.3.4 8.3.4. Enzyme Design

### v2

@@ 7. Databases for Protein Representation Learning
Table 2. Summary of widely used databases in Protein Representation Learning (PRL), categorized by sequence, structure, and function.
Name Data Type Scale Description
UniProtKB (boutet2007uniprotkb, ; EDITtoTrEMBL, ) Sequence, Function $\sim$252M Comprises Swiss-Prot (manually curated) and TrEMBL (automatically annotated); widely used in pretraining and function annotation.
UniRef (suzek2007uniref, ) Sequence $\sim$250M Clusters UniProt sequences at different identity thresholds to reduce redundancy and improve training efficiency.
Pfam (mistry2021pfam, ) Sequence $\sim$22K Defines protein families and domains using MSAs and profile Hidden Markov Models to support motif-level annotation.
@@
MGnify (richardson2023mgnify, ) Sequence $\sim$2.4B Provides non-redundant protein sequences predicted from metagenomic environmental samples to increase diversity.
@@
PDB (berman2000protein, ) Structure $\sim$231K Repository of experimentally determined protein structures via X-ray crystallography, NMR, and cryo-EM.
@@
AlphaFold DB (varadi2024alphafold, ) Structure $\sim$214M Stores high-confidence predicted protein structures from AlphaFold to improve structural coverage.
ESMAtlas (lin2023evolutionary, ) Structure $\sim$772M Contains predicted structures from ESMFold, including many novel and uncharacterized proteins.
@@
Gene Ontology (gene2023gene, ) Function $\sim$45M Provides structured functional annotations across biological processes, molecular functions, and cellular components.
8. Applications
@@
Protein Representation Learning (PRL) has revolutionized biological research and biomedical innovation, providing data-driven solutions to understand and engineering proteins. By learning meaningful representations of protein sequences, structures, and interactions, PRL can enable a wide range of applications, from predicting fundamental protein properties to advancing drug discovery. As illustrated in Fig. 4, PRL applications can be broadly categorized into four main domains: prediction of protein properties, prediction of protein structure, design and optimization of proteins and drug discovery.
@@
Figure 4. Overview of key applications of Protein Representation Learning (PRL). PRL enables advancements in multiple domains, including protein property prediction, structure prediction, protein design and optimization, and drug discovery.
8.1. Protein Property Prediction
@@
Benchmarks for evaluating protein property prediction models are categorized according to the type of property that is being assessed. For individual protein properties, Tasks Assessing Protein Embeddings (TAPE) (rao2019evaluating, ) provides a standardized evaluation framework, including fluorescence prediction and protease stability assessment. Another widely used benchmark, UniProt (boutet2007uniprotkb, ), serves as a comprehensive database of protein sequences, functional annotations, and metadata from various organisms, offering a valuable resource for assessing sequence-based property predictions in various protein families and functional contexts. In contrast, interaction-based properties, such as protein-ligand and protein-protein interactions, require different benchmarks that incorporate structural and binding information. Prominent datasets in this category include PDBBind (PDBbind, ), Leak Proof PDBBind (li2024leakproofpdbbindreorganized, ), Davis (Davis, ), KIBA (KIBA, ), PLINDER (Durairaj2024.07.17.603955, ) and PINDER (kovtun2024pinder, ), which evaluate models based on binding affinity predictions. These datasets serve as widely recognized standards, enabling consistent model comparison and performance assessment across different protein property prediction tasks.
@@
With the remarkable advancements in Protein Language Models (PLMs), most current research leverages these models to extract meaningful representations of protein sequences, applying them to a wide range of property prediction tasks. These models have been successfully employed for both individual protein properties (yang2018learned, ; yu2023enzyme, ; buton2023predicting, ) as well as interaction-based properties (yuan2025mutualdata, ; yuan2023dg, ; ha2025lantern, ). Beyond sequence-based representations, many efforts have explored the integration of multiple modalities, such as structural, physicochemical, and functional data, to further enhance protein representations beyond what PLM alone can achieve (wang2022lm, ; nguyen2024multimodal, ; ngo2024multimodal, ; luo2024enhancing, ). By incorporating additional biological information, these approaches aim to provide a more comprehensive view of protein function and interactions, particularly in tasks where spatial and contextual features play a significant role, such as protein-ligand binding and protein-protein interactions.
8.2. Protein Structure Prediction
@@
Protein structure prediction, also known as protein folding, aims to determine the three-dimensional conformation of a protein from its amino acid sequence. Since protein structure dictates function, accurately predicting spatial arrangements is essential for understanding molecular interactions, drug binding, and enzymatic activity. Traditionally, protein structures have been determined using experimental techniques such as X-ray crystallography (https://doi.org/10.1111/febs.12796, ) and cryo-electron microscopy (Cryo-EM) (BAI201549, ), which provide high-resolution structural insights. However, these methods are often time-consuming, costly, and limited by challenges such as crystallization feasibility and sample preparation (venien2017cryo, ). Advances in computational structure prediction have transformed the field by providing faster, scalable alternatives that complement experimental approaches, broadening the scope of protein engineering and biomedical research.
@@
To evaluate the accuracy of structure prediction models, several benchmarking platforms have been established. The Critical Assessment of Techniques for Protein Structure Prediction (CASP) (kryshtafovych2021critical, ), a biennial blind assessment, is widely regarded as the gold standard for evaluating protein folding models. In addition, Continuous Automated Model EvaluatiOn (CAMEO) (robin2021continuous, ) provides real-time assessments by continuously comparing predictions against newly released experimental structures. These benchmarks are instrumental in validating and refining computational models, ensuring their reliability and applicability in structural biology.
@@
Protein structure prediction models can be broadly categorized into MSA-based and MSA-free approaches. MSA-based methods leverage multiple sequence alignments to capture evolutionary relationships, often resulting in higher accuracy for proteins with rich evolutionary data. Notable models in this category include AlphaFold (jumper2021highly, ), RosettaFold (baek2021accurate, ), OpenFold (ahdritz2024openfold, ), and MSAGPT (chen2024msagpt, ). In contrast, MSA-free methods predict structures directly from single sequences without relying on evolutionary context, making them more efficient and applicable to proteins with limited sequence homology. Examples of MSA-free models include ESMFold (lin2023evolutionary, ), OmegaFold (wu2022high, ), and HelixFold-Single (fang2023method, ). While these models have greatly improved protein structure prediction, they are primarily limited to static conformations, as they are trained on crystallographic and experimentally resolved structures. However, proteins are inherently dynamic and undergo conformational changes essential to their function. To address this limitation, recent advancements have introduced generative models capable of predicting conformational ensembles, such as BioEmu (lewis2024scalable, ).
8.3. Protein Design and Optimization
@@
Protein design and optimization involve strategies for engineering proteins with specific structural and functional properties. Broadly, these strategies fall into two main categories: sequence-based design, which optimizes amino acid sequences to achieve target properties, and structure-based design, which leverages protein structural features either by generating backbones and applying inverse folding to identify compatible sequences or by co-designing sequence and structure to enhance functional and structural outcomes. These approaches have been applied to various protein types, including individual proteins with optimized functionality, ligand-binding proteins for molecular interactions, antibodies for therapeutic applications, and enzymes for improved catalytic efficiency in industrial and biomedical settings.
8.3.1. Individual Protein Property Optimization
@@
Individual protein property optimization focuses on enhancing the intrinsic characteristics of a protein, such as stability, activity, specificity, and catalytic efficiency, without explicitly considering interactions with other molecules. Traditional approaches, such as directed evolution with random mutagenesis (doi:10.1021/ar960017f, ; TRACEWELL20093, ; doi:10.1073/pnas.1215251110, ), iteratively introduce mutations and select improved variants. While these methods have been widely used and have led to many successful optimizations, they often require extensive experimental screening. Computational strategies have been developed to facilitate a more systematic exploration of sequence space, offering alternatives that reduce experimental reliance while aiming to improve efficiency in protein engineering.
@@
To evaluate optimization strategies, several benchmark datasets have been established for sequence-based optimization, each representing distinct challenges between protein targets, originating organisms, sequence lengths, dataset sizes, and fitness objectives. Notable benchmarks include Green Fluorescent Protein (avGFP) (sarkisyan2016local, ), Adeno-Associated Viruses (AAV) (bryant2021deep, ), TEM-1 $\beta$-Lactamase (TEM) (firnberg2014comprehensive, ), Ubiquitination Factor Ube4b (E4B) (starita2013activity, ), Aliphatic Amide Hydrolase (AMIE) (Wrenbeck2017, ), Levoglucosan Kinase (LGK) (Klesmith2015, ), Poly(A)-binding Protein (Pab1) (Melamed_2013, ), and SUMO E2 Conjugase (UBE2I) (weile2017framework, ). These datasets enable a standardized assessment of different protein optimization methods.
@@
Earlier approaches focused on optimizing proteins in discrete sequence space. These methods explored mutation-based search strategies, reinforcement learning, and PLMs to propose beneficial sequence modifications (emami2023plug, ; tran2024protein, ). Building upon this foundation, recent advancements in PRL have introduced computational frameworks capable of mapping protein sequences into fitness landscapes. These models aim to capture sequence-function relationships and facilitate direct sequence optimization based on learned representations. Unlike experimental screening, PRL-based approaches rely on large-scale protein datasets to infer meaningful sequence features and predict fitness variations. A common strategy in fitness landscape modeling integrates sequence and fitness information within a machine learning framework. One such approach involves jointly training an autoencoder with a prediction network, where an encoder extracts sequence features, a decoder reconstructs sequences, and a separate model predicts fitness values from the latent space (castro2022transformer, ; 10.1088/2632-2153/adc2e2, ). Protein Language Models (PLMs) are often used as encoders to capture contextual sequence representations, and the prediction network guides latent space organization based on fitness variations.
@@
Other studies have explored methods to improve fitness landscape modeling through smoothing techniques, aiming to enhance search efficiency in sequence space. Graph-based smoothing techniques have been introduced to incorporate protein similarity graphs, promoting structural continuity in the fitness landscape (kirjner2023improving, ; tran2024groot, ). These methods aim to reduce abrupt fitness changes caused by minor sequence variations, improving the robustness of optimization algorithms. In general, while experimental and computational strategies each offer advantages and limitations, ongoing research continues to explore ways to enhance protein optimization efficiency while ensuring the reliability and interpretability of computational predictions.
8.3.2. Ligand-Binding Protein Design
@@
The design of ligand-binding proteins involves engineering proteins that interact with small molecules such as drugs, metabolites, or signaling compounds. These interactions are central to biological processes, making the design of ligand-binding proteins essential for applications in biosensors, precision medicine, and enzyme-mediated synthesis. Achieving high specificity and affinity for target ligands requires a detailed understanding of the binding site geometry, molecular interactions, and binding thermodynamics, which computational methods have been increasingly employed to model and optimize.
@@
Several benchmark datasets support the evaluation of ligand-binding protein design strategies by providing experimentally validated protein-ligand complexes. PDBbind (PDBbind, ) offers a curated collection of protein-ligand structures with binding affinity measurements, serving as a reference for affinity prediction and binding site modeling. CSAR (Community Structure Activity Resource) (dunbar2013csar, ) provides high-resolution crystallographic data with a diverse range of ligand complexes, offering a structured framework for evaluating the specificity of the binding across various protein targets. These datasets are widely used to assess structure-based and sequence-based ligand-binding design approaches.
@@
Computational strategies for the design of ligand-binding proteins typically leverage structural insights from native protein-ligand complexes to refine side chain interactions and backbone conformations, with the aim of improving binding affinity (polizzi2020defined, ; stark2023harmonic, ; watson2023novo, ; dauparas2023atomic, ). Many approaches follow a structure-first paradigm, in which protein backbones are generated first, followed by inverse folding techniques (hsu2022learning, ; gao2023pifold, ) to identify sequences that can adopt predefined structural conformations. This prioritization of structure prediction before sequence determination allows the design of proteins that conform to specific ligand-binding geometries and can also aid in addressing cases where binding sites are not well defined.
@@
Despite progress, challenges remain, particularly in the structural validation of designed binding modes. Although computational coupling has been applied to generate novel binders by modifying scaffolds and loop geometries, the accuracy and stability of the predicted interactions often require high-resolution experimental confirmation. Furthermore, many design strategies emphasize a limited set of hotspot residues for scaffold placement, which can restrict the exploration of various interaction modes, particularly for targets lacking well-defined pockets or binding clefts (Pocketminer, ; Gagliardi2023-do, ). To address these limitations, alternative approaches have investigated the design of ligand-binding proteins based only on sequence information, bypassing the reliance on structural data (nguyen2024proteinrediff, ). These models leverage Protein Language Models (PLMs) to extract functional sequence motifs associated with ligand binding, allowing the design of proteins with desired binding properties even when structural information is unavailable or difficult to obtain.
8.3.3. Antibody Design
@@
Antibodies are specialized immune proteins that recognize and bind to specific antigens, such as pathogens, toxins, or diseased cells, marking them for neutralization or destruction. As central components of adaptive immunity, antibodies exhibit high specificity and strong binding affinity, allowing precise immune responses. Their function is primarily dictated by variable regions, particularly Complementarity Determining Regions (CDRs), which mediate antigen recognition. Antibody design focuses on the engineering of antibodies with improved affinity, specificity, and stability, with broad therapeutic applications in cancer, autoimmune disorders, and infectious diseases.
@@
Benchmark datasets play a critical role in evaluating sequence-based and structure-based antibody design. OAS (Observed Antibody Space) (olsen2022observed, ) provides more than 2 billion immune repertoire sequences from various immune states, species, and individuals, supporting sequence-based antibody modeling. SAbDab (Structural Antibody Database) (dunbar2014sabdab, ) compiles annotated antibody structures from the PDB, including affinity data and sequence annotations, serving as a key resource for structure-based antibody modeling. AbBind (sirin2016ab, ) includes 1,101 antibody-antigen mutants in 32 complexes with experimentally measured binding free energy changes, allowing for the study of affinity variations upon mutation. These datasets support the evaluation and development of antibody design approaches, ensuring standardized model assessment across different strategies.
@@
Computational strategies for structure-based antibody design focus primarily on CDRH3 loops, which are highly variable and critical for antigen binding. Two main methodologies have emerged: (i) 3D backbone generation, which designs structurally realistic CDRH3 loops for stable antigen recognition (eguchi2022ig, ), and (ii) prediction of binding affinity, which models how sequence modifications impact structural stability and antigen binding energy (shan2022deep, ). While structure-based approaches have been widely used, a major challenge is the limited availability of high-resolution antibody structures, which constrains direct structural modeling. To address this, sequence-based models leverage Protein Language Models (PLMs) trained on large-scale immune repertoires to learn patterns associated with high-affinity antibodies. These methods enable de novo antibody generation and affinity maturation without requiring structural information (cohen2023epitopespecific, ). Furthermore, sequence-based antibody design can be framed as an optimization problem, where binding affinity serves as the fitness metric, guiding the search for improved variants, similar to optimization strategies used for individual protein properties (wu2025a, ; lin2025dyab, ). A promising direction in antibody design is the integration of sequence and structural information, leading to sequence-structure co-design approaches (jin2022iterative, ; luo2022antigen, ; morcillo2023guiding, ; martinkus2023abdiffuser, ; zhou2024antigen, ). By combining both modalities, these models aim to capture the relationships between sequence variability and structural adaptation, facilitating the precise engineering of antibody specificity, stability, and affinity. This hybrid framework expands the design space beyond naturally occurring immune repertoires, potentially accelerating antibody discovery and improving antigen recognition strategies.
8.3.4. Enzyme Design